If a contractor charges 30 for 2hours work and charges 94 for 4 hours of work what would he charge for 8 hours of work?
Answer by tutorcecilia(2152)

Since the amount of money charged increases with the amount of time spent working, there is a linear relationship between time and money. One type of equation you could use is the slope-intercept equation of a line (y=mx+b)
.
First find the slope of the line by using the two points that are given:
(2, 30) and (4, 94)
.
Slope =m=(y-y2)/(x-x2) [Formula for the slope of a line]
.
m=(94-30)/(4-2)=64/2=32 [Plug-in the values of the two points]
.
y=mx+b
30=(32)(2)+b [Plug-in the values for the slope and one point]
30=65+b [Solve for b]
-65+30=b
-34=b
.
y=mx+b [Plug-in the values for the slope and the y-intercept (b)]
y=32x-34 [Equation of the line]
.
To find any point on the line, plug-in the known variable and solve for the unknown variable:
.
Find (8, y)
y=(32)(8)-34
y=222
So, when 8 hours are worked, the charge is $222